What's the value of x if f(x)=9
f(x)=5x–1

Answer by jim_thompson5910(35256)

f(x)=5x–1

9 = 5x–1 ... replace f(x) with 9 (since we're told that f(x) = 9)

9+1 = 5x–1+1

10 = 5x

5x = 10

5x/5 = 10/5

x = 2